A visitor to a lighthouse wishes to determine
the height of the tower. She ties a spool
of thread to a small rock to make a simple
pendulum, which she hangs down the center
of a spiral staircase of the tower. The period
of oscillation is 10.8 s.
The acceleration of gravity is 9.8 m/s2 .
What is the height of the tower?
Answer in units of m.
1
Expert's answer
2011-01-18T02:46:28-0500
The formula for the period of oscillations can be written as
T = 2 π √(L/g)

L = gT2/4 π2
L = 9.8*10.82 / 4π2 = 28.95 m.